### v1

@@ 3 Properties of Fejér* monotonicity
$\beta_{k+1}=\sqrt{\alpha_{k}^{2}-2\alpha_{k}+\beta_{k}^{2}}\leq\sqrt{\alpha_{k}^{2}-2\alpha_{k}\lambda+\beta_{k}^{2}}\eqqcolon\gamma_{k},$
and $\gamma_{k}$ is such that $z^{k}\coloneqq(0,\gamma_{k})$ belongs to the circumference centered in $w$ and passing through $x^{k}$. In other words, $z^{k}$ complies with $\|z^{k}-w\|=\|x^{k}-w\|$. Thus,
$\|x^{k+1}-w\|=\sqrt{\beta_{k+1}^{2}+\lambda^{2}}\leq\sqrt{\gamma_{k}^{2}+\lambda^{2}}=\|x^{k}-w\|.$
On the other hand, for $w^{\ell}=(1/2^{\ell},0)\in{M}$ with $\ell\geq 0$, consider $N(w^{\ell})=2\ell$. For all even $k\geq N(w^{\ell})$, we have $\alpha_{k}=0$, $\alpha_{k+1}\leq 1/2^{\ell}$ and $\beta_{k+1}=\beta_{k}$. So, $\|x^{k+1}-w^{\ell}\|\leq\|x^{k}-w^{\ell}\|$ and the sequence is Fejér* monotone w.r.t. $M$.
Note that $\operatorname{conv}(M)=(0,1]\times\{0\}$ and also it holds that $\bar{w}=(0,0)\in\operatorname{cl}\left({\operatorname{conv}(M)}\right)$. We show now that $(x^{k})_{k\in\mathds{N}}$ is not Fejér* monotone w.r.t. $\{(0,0)\}$. Indeed, for $k=2\ell$, $\ell\in\mathds{N}$, we have $x^{2\ell}=(0,\beta_{2\ell})$ and $x^{{2\ell}+1}=(1/2^{\ell},\beta_{{2\ell}+1})=(1/2^{\ell},\beta_{2\ell})$. Thus,
$\|x^{{2\ell}+1}-\bar{w}\|^{2}=\dfrac{1}{2^{2\ell}}+\beta_{2\ell}^{2}>\beta_{2\ell}^{2}=\|x^{2\ell}-\bar{w}\|^{2}.$
Finally, just for the sake of completion, we show that sequence $(x^{k})_{k\in\mathds{N}}$ converges to $x^{*}\coloneqq(0,\sqrt{4/3})$. In fact, for $k=2\ell$, $\ell\in\mathds{N}$, we have from Example 1 that
@@
Clearly, we have $\left\lVert x^{2\ell+1}-x^{2\ell}\right\rVert\to 0$ when $\ell\to+\infty$, so $x^{k}\to x^{*}$, as desired. ∎
Remark 1
@@
One might argue that the set $M$ in Example 1 is somehow special, for $M$ has empty interior. However, it is possible to prove the same result for an extended example where the underlying set has nonempty interior, say, for instance, $M\coloneqq\{(w^{\prime},w^{\prime\prime})\in\mathds{R}^{2}\mid 0<w^{\prime}\leq 1,-1<w^{\prime\prime}\leq 0\}$.
@@
In the next proposition the closure of the convex hull of the underlying set plays a role. Even though a Fejér* monotone sequence w.r.t. $M\subset\mathcal{H}$ is not Fejér* monotone w.r.t. $\operatorname{cl}\left({\operatorname{conv}(M)}\right)$, one can characterize Fejér* monotonicity in terms of the convergence of scalar sequences and inner product sequences to the closure of the convex hull. This result can be useful for the analysis regarding weak convergence of optimization algorithms that generate Fejér* monotone sequences.
Proposition 2
@@
Let $(x^{k})_{k\in\mathds{N}}\subset\mathcal{H}$ be a Fejér* monotone w.r.t. a nonempty set $M$ in $\mathcal{H}$. Then,
@@
    for every $\bar{x}\in\operatorname{cl}\left({\operatorname{conv}(M)}\right)$ the scalar sequence $(\left\lVert x^{k}-\bar{x}\right\rVert)_{k\in\mathds{N}}$ converges;
@@
Proof
Regarding item (i), recall that Proposition 1(iii) implies that $(x^{k})_{k\in\mathds{N}}$ is Fejér* monotone w.r.t. $\operatorname{conv}(M)$. Thus, for any $w\in\operatorname{conv}(M)$, Proposition 1(ii) yields that the scalar sequence $(\left\lVert x^{k}-w\right\rVert)_{k\in\mathds{N}}$ converges.
Now, take $\bar{x}\in\operatorname{cl}\left({\operatorname{conv}(M)}\right)$, and consider a sequence $(w^{\ell})_{k\in\mathds{N}}\subset\operatorname{conv}(M)$ such that $w^{\ell}\rightarrow\bar{x}$. Using the triangle inequality, we have, for all $\ell\in\mathds{N}$,
@@
$-\left\|w^{\ell}-\bar{x}\right\|$ $\leq\liminf_{k}\left\|x^{k}-\bar{x}\right\|-\lim_{k}\left\|x^{k}-w^{\ell}\right\|$ (7)
$\leq\limsup_{k}\left\lVert x^{k}-\bar{x}\right\rVert-\lim_{k}\|x^{k}-w^{\ell}\|\leq\left\|w^{\ell}-\bar{x}\right\|,$ (8)
because, as aforementioned, item Proposition 1(ii) says that, for every $\ell\in\mathds{N}$, $\lim_{k}\left\|w^{\ell}-x^{k}\right\|$ exists. Now, as $\ell$ goes to infinity, we have $\left\|w^{\ell}-\bar{x}\right\|\to 0$, and we conclude that
$\liminf_{k}\left\|x^{k}-\bar{x}\right\|=\limsup_{k}\left\|x^{k}-\bar{x}\right\|=\lim_{k}\left\|x^{k}-\bar{x}\right\|,$
@@
Now, item (i) implies that the sequences $(\left\|x^{k}-\bar{x}_{1}\right\|)_{k\in\mathds{N}}$ and $(\left\|x^{k}-\bar{x}_{2}\right\|)_{k\in\mathds{N}}$ converge. Therefore, the sequence $\left(\left\langle\bar{x}_{1}-\bar{x}_{2},x^{k}\right\rangle\right)_{k\in\mathds{N}}$ converges. ∎
@@ 4 Fejér* and quasi-Fejér sequences
In this section we discuss Fejér* monotonicity relation with quasi-Fejér monotone sequences (as addressed in Combettes:2001a ). First, we establish the following notation: $\ell_{+}$ is the set of real sequences with nonnegative entries, and $\ell^{1}$ is the set of real sequences with finite $\ell^{1}$-norm.
Definition 2 (quasi-Fejér monotonicity (Combettes:2001a, , Def. 1.1))
@@
    $\left\|x^{k+1}-x\right\|\leq\left\|x^{k}-x\right\|+\varepsilon_{k};$
@@
    $\left\|x^{k+1}-x\right\|^{2}\leq\left\|x^{k}-x\right\|^{2}+\varepsilon_{k};$
@@
    $\left\|x^{k+1}-x\right\|^{2}\leq\left\|x^{k}-x\right\|^{2}+\varepsilon_{k}.$
Sure enough, Fejér sequences are quasi-Fejér of Type I sequences. Also, one can clearly see that Type II implies Type III. Moreover, it was established in (Combettes:2001a, , Prop. 3.2) that if $M$ is bounded, then Type I implies Type II.
Quasi-Fejér of Type III sequences are the most general ones. However, due to this generality, they are hardly used for the analysis of optimization algorithms. In fact, the most common type of quasi-Fejér sequences used in the literature are quasi-Fejér of Type II sequences; see, for instance, Alber:1998 ; Iusem:1994 .
The next theorem shows that Fejér* monotonicity sequences are also quasi-Fejér of Type III.
Theorem 4.1
Let $(x^{k})_{k\in\mathds{N}}\subset\mathcal{H}$ be a Fejér* monotone sequence w.r.t. a nonempty set $M$ in $\mathcal{H}$. Then, $(x^{k})_{k\in\mathds{N}}$ is a quasi-Fejér of Type III sequence w.r.t. $M$.
Proof
Suppose that $(x^{k})_{k\in\mathds{N}}$ is Fejér* monotone w.r.t. $M$. Then, in view of Definition 1, for any $x\in M$, there exists $N(x)$ such that, for all $k\geq N(x)$, it holds that
$\left\|x^{k+1}-x\right\|^{2}\leq\left\|x^{k}-x\right\|^{2}.$
@@
$\left\|x^{k+1}-x\right\|>\left\|x^{k}-x\right\|,$
@@
$\varepsilon_{k}\coloneqq\begin{cases*}\left\|x^{k+1}-x\right\|^{2}-\left\|x^{k}-x\right\|^{2},&\text{if $\left\|x^{k+1}-x\right\|>\left\|x^{k}-x\right\|$},\\ 0,&\text{otherwise}.\end{cases*}$
Thus, it holds true that sequence $(\varepsilon_{k})_{k\in\mathds{N}}\in\ell_{+}\cap\ell^{1}$ and therefore sequence $(x^{k})_{k\in\mathds{N}}$ is quasi-Fejér of Type III w.r.t. $M$. ∎
We note that the concepts of Fejér* and quasi-Fejér of Type III sequences clearly are not equivalent, with the latter being more general than the former.
Actually, if $M$ is a singleton, then any Fejér* monotone sequence is quasi-Fejér of Type II, using the same arguments used in the proof of Theorem 4.1. Of course, in this case one can easily build quasi-Fejér sequences of Type II that are not Fejér*. Nevertheless, if a sequence is quasi-Fejér of Type II w.r.t. a general $M$, then it is Type II with respect to $\operatorname{cl}\left({\operatorname{conv}(M)}\right)$ (see (Combettes:2001a, , Prop. 3.4)). Hence, in view of Example 1, we have a Fejér* sequence that it is not quasi-Fejér of Type II monotone. Note that we can construct a non-singleton set $M$, for instance, $M\coloneqq[-1,0]\times\{0\}\subset\mathds{R}^{2}$, such that the same sequence $(x^{k})_{k\in\mathds{N}}$ of Example 1 is quasi-Fejér of Type II, but not Fejér* monotone, w.r.t. $M$. More than not being equivalent, we have just seen that the concepts of Fejér* and quasi-Fejér of Type II are not contained in each other, that is, those ideas are independent.
The previous theorem reveals a particular quasi-Fejér of type III sequence that is related to optimization algorithms, through Fejér* monotonicity Behling:2024c . Thus, we end our discussion by presenting some characterizations of Fejér* monotonicity in terms of the weak convergence of the underlying sequence. These statements are inspired by Combettes:2001a and could be derived from the results therein, since Fejér* monotonicity implies quasi-Fejér of type III monotonicity. However, we present them here, for the sake of completeness. Moreover, the provided proofs are simpler and more clear for the reader, in the context of Fejér* monotonicity.
From now on, for a sequence $(x^{k})_{k\in\mathds{N}}\subset\mathcal{H}$, we denote by $\mathfrak{W}\left(x^{k}\right)_{k\in\mathds{N}}$ the set of its weak cluster points and by $\mathfrak{S}\left(x^{k}\right)_{k\in\mathds{N}}$ the set of its strong cluster points. First, we give portrayals of Fejér* monotonicity in terms of the weak convergence of the sequence.
Proposition 3
Let $\left(x^{k}\right)_{k\in\mathds{N}}$ be a Fejér* sequence with respect to $M\subset\mathcal{H}$. Then,
@@
    $M\subset\left\{y\in\mathcal{H}\mid\left\langle y,w^{\prime}-w^{\prime\prime}\right\rangle=\alpha\right\};$
@@
Proof
Item (i): the statement is valid since, due to Proposition 1(i), sequence $\left(x^{k}\right)_{k\in\mathds{N}}$ is bounded.
@@
$\left\|x^{k}-y\right\|^{2}-\|y\|^{2}=\left\|x^{k}\right\|^{2}-2\left\langle y,x^{k}\right\rangle.$
Moreover, it follows from Proposition 1(ii) that $\lim\left(\left\|x^{k}-y\right\|^{2}-\|y\|^{2}\right)$ is well-defined.
Now, given two points $w^{\prime},w^{\prime\prime}\in\mathfrak{W}\left(x^{k}\right)_{k\in\mathds{N}}$, such that $x^{j_{k}}\rightharpoonup w^{\prime}$ and $x^{\ell_{k}}\rightharpoonup w^{\prime\prime}$, and considering Section 4 we have
$\lim\left\|x^{j_{k}}\right\|^{2}-2\left\langle y,w^{\prime}\right\rangle=\lim\left\|x^{\ell_{k}}\right\|^{2}-2\left\langle y,w^{\prime\prime}\right\rangle,$

### v2

@@ 3 Properties of Fejér* monotonicity
$\beta_{k+1}=\sqrt{\alpha_{k}^{2}-2\alpha_{k}+\beta_{k}^{2}}\leq\sqrt{\alpha_{k}^{2}-2\alpha_{k}\lambda+\beta_{k}^{2}}\eqqcolon\gamma_{k},$ (13)
and $\gamma_{k}$ is such that $z^{k}\coloneqq(0,\gamma_{k})$ belongs to the circle centered in $w$ and passing through $x^{k}$. A simple calculation shows that $\|z^{k}-w\|=\|x^{k}-w\|$. Thus,
$\|x^{k+1}-w\|=\sqrt{\beta_{k+1}^{2}+\lambda^{2}}\leq\sqrt{\gamma_{k}^{2}+\lambda^{2}}=\|x^{k}-w\|.$ (14)
On the other hand, for $w^{\ell}=(1/2^{\ell},0)\in{M}$ with $\ell\geq 0$, consider $N(w^{\ell})=2\ell$. For all even $k\geq N(w^{\ell})$, we have $\alpha_{k}=0$, $\alpha_{k+1}\leq 1/2^{\ell}$ and $\beta_{k+1}=\beta_{k}$. So, $\|x^{k+1}-w^{\ell}\|\leq\|x^{k}-w^{\ell}\|$ and the sequence is Fejér* monotone with respect to $M$.
Note that $\operatorname{conv}(M)=(0,1]\times\{0\}$ and also it holds that $\bar{w}=(0,0)\in\operatorname{cl}\left({\operatorname{conv}(M)}\right)$. We show now that $(x^{k})_{k\in\mathds{N}}$ is not Fejér* monotone with respect to $\{(0,0)\}$. Indeed, for $k=2\ell$, $\ell\in\mathds{N}$, we have $x^{2\ell}=(0,\beta_{2\ell})$ and $x^{{2\ell}+1}=(1/2^{\ell},\beta_{{2\ell}+1})=(1/2^{\ell},\beta_{2\ell})$. Thus,
$\|x^{{2\ell}+1}-\bar{w}\|^{2}=\dfrac{1}{2^{2\ell}}+\beta_{2\ell}^{2}>\beta_{2\ell}^{2}=\|x^{2\ell}-\bar{w}\|^{2}.$ (15)
Finally, just for the sake of completion, we show that sequence $(x^{k})_{k\in\mathds{N}}$ converges to $x^{*}\coloneqq(0,\sqrt{4/3})$. In fact, for $k=2\ell$, $\ell\in\mathds{N}$, we have from Eq.˜13 that
@@
Clearly, we have $\left\lVert x^{2\ell+1}-x^{2\ell}\right\rVert\to 0$ when $\ell\to+\infty$, so $x^{k}\to x^{*}$, as desired.
Remark 3.4.
@@
One might argue that the set $M$ as defined in Example˜3.3 is somehow special, for $M$ has empty interior. However, it is possible to prove the same result for an extended example where the underlying set has nonempty interior, say, for instance, the square
$S\coloneqq\{(w^{\prime},w^{\prime\prime})\in\mathds{R}^{2}\mid 0<w^{\prime}\leq 1,-1<w^{\prime\prime}\leq 0\},$ (18)
instead of $M$. Note that $\operatorname{conv}(M)\subsetneq(0,1]\times(-1,0]=S$.
We present now two counter-examples showing convex combination of Fejér* monotone sequences may fail to be Fejér* monotone, as well as the fact that composition of Fejér* monotone sequences may fail to be Fejér* monotone.
We start with the latter. By composition of Fejér* monotone sequences we mean the following. Let $T_{1},T_{2}:\mathcal{H}\to\mathcal{H}$ be two operators that generate Fejér* sequences with respect to a target set $C\subset\mathcal{H}$, that is, sequences $(x^{k}_{i})_{k\in\mathds{N}}$ defined by $x^{k+1}_{i}=T_{i}(x^{k}_{i})$ with starting point $x^{0}_{i}\in\mathcal{H}$, for $i=1,2$, that are Fejér* monotone with respect to $C$. Consider the operator $T=T_{1}\circ T_{2}$ and the sequence $(x^{k})_{k\in\mathds{N}}$ defined by $x^{k+1}=T(x^{k})=T_{1}(T_{2}(x^{k}))$ with starting point $x^{0}\in\mathcal{H}$. The question is whether the sequence $(x^{k})_{k\in\mathds{N}}$ is Fejér* monotone with respect to $C$. The next example shows that this is not necessarily the case.
Example 3.5.
Let $\mathcal{H}=\mathds{R}$ and let the target set be $C=\{0\}$. We define three auxiliary sequences for $k\in\mathds{N}$:
$u^{k}=4^{-k},\quad v^{k}=2\cdot 4^{-k},\quad w^{k}=-4^{-(k+1)}.$ (19)
We define the operators $T_{1},T_{2}:\mathds{R}\to\mathds{R}$ by specifying their action in the following manner:
$T_{1}(x)$ $=\begin{cases}v^{k}&\text{if }x=u^{k},\\ 0&\text{if }x=v^{k},\\ u^{k+1}&\text{if }x=w^{k},\\ 0&\text{otherwise},\end{cases}$ $T_{2}(x)$ $=\begin{cases}u^{k}&\text{if }x=u^{k},\\ w^{k}&\text{if }x=v^{k},\\ 0&\text{if }x=w^{k},\\ 0&\text{otherwise}.\end{cases}$ (20)
We claim now that both $T_{1}$ and $T_{2}$ generate Fejér* sequences with respect to the target set $C=\{0\}$. Indeed, for $T_{1}$, a trajectory starting at $u^{k}$ moves $u^{k}\to v^{k}\to 0$. Although $|v^{k}|>|u^{k}|$ (an expansion), the sequence hits the target in 2 steps. A trajectory starting at $w^{k}$ moves $w^{k}\to u^{k+1}\to v^{k+1}\to 0$. The remaining trajectories go directly to the target. In all cases, the sequence enters the target set after finitely many steps, so the tail is monotone (constant zero). Thus, $T_{1}$ generates Fejér* sequences.
For $T_{2}$, a trajectory starting at $v^{k}$ moves $v^{k}\to w^{k}\to 0$. Since $|w^{k}|<|v^{k}|$, this is strictly contractive. Trajectories starting at $u^{k}$ stay fixed at $u^{k}$ (monotone). The remaining trajectories go directly to the target. Thus, $T_{2}$ also yields Fejér* sequences.
Consider now the operator $T=T_{1}\circ T_{2}$ and the sequence $(x^{k})_{k\in\mathds{N}}$ defined by $x^{k+1}=T(x^{k})$ with starting point $x^{0}=u^{0}=1$. Then, the sequence $(x^{k})_{k\in\mathds{N}}$ evolves as follows:
$u^{0}\to v^{0}\to u^{1}\to v^{1}\to u^{2}\to v^{2}\to\dots$ (21)
Explicitly, we analyse the transitions.
From $u^{k}$ to $v^{k}$, we have $T_{2}(u^{k})=u^{k}$, and then $T_{1}(u^{k})=v^{k}$. Thus, if $x^{k}=u^{k}$, then $x^{k+1}=v^{k}$. Here, the distance to the target $C$ changes from $|u^{k}|$ to $|v^{k}|=2|u^{k}|$, yielding $|x^{k+1}-0\ >|x^{k}-0|$, a violation of Fejér* monotonicity.
Now, from $v^{k}$ to $u^{k+1}$ we get $T_{2}(v^{k})=w^{k}$, and then $T_{1}(w^{k})=u^{k+1}$. Tehrefore, if $x^{k+1}=v^{k}$, then $x^{k+2}=u^{k+1}$. Now, the distance changes from $|v^{k}|$ to $|u^{k+1}|=\frac{1}{8}|v^{k}|$, which corresponds to a contraction.
This indicates that the sequence generated by $T$ oscillates indefinitely between expansion and contraction. Since the distance strictly increases at every step mapping $u^{k}\to v^{k}$ (which occurs infinitely often), there exists no index $K$ such that the tail of the sequence is monotone non-increasing. Therefore, the composition $T$ does not generate a Fejér* sequence.
We now address the case of convex combinations of Fejér* monotone sequences.
Example 3.6.
Let $\mathcal{H}=\mathds{R}$ and $C=\{0\}$. Let $(x^{k})_{k\in\mathds{N}}$ and $(y^{k})_{k\in\mathds{N}}$ be two sequences defined by: $x^{0}=2,x^{1}=1,x^{k}=2/k$ and $y^{0}=-2,y^{1}=-1.8,y^{k}=-2+2\cdot 10^{-k}$, for $k\geq 1$. Both sequences are Fejér monotone with respect to $C$, since both converge to zero. However, consider their convex combination $w^{k}\coloneqq\frac{1}{2}x^{k}+\frac{1}{2}y^{k}$. We have $w^{0}=0$, which implies $\left\lvert w^{0}-0\right\rvert=0$. At the next step, $w^{1}=\frac{1}{2}(1-1.8)=-0.4$, yielding $\left\lvert w^{1}-0\right\rvert=0.4$. Note that $\left\lvert w^{1}-0\right\rvert>\left\lvert w^{0}-0\right\rvert$. Indeed, $w^{k}=-1+1/k+10^{-k}$, for $k\geq 1$, which is decreasing sequence converging to 1-1, yet with $\left\lvert w^{k+1}\right\rvert=\left\lvert w^{k+1}-0\right\rvert>\left\lvert w^{k}-0\right\rvert=\left\lvert w^{k}\right\rvert$. Thus, the sequence $(w^{k})_{k\in\mathds{N}}$ fails to be Fejér monotone with respect to $C=\{0\}$, despite being a convex combination of two Fejér sequences (converging to zero). Since Fejér monotone sequences are Fejér* sequences, this example also shows that convex combinations of Fejér* sequences are not necessarily Fejér* sequences.
The next remark gives a sufficient condition for convex combinations of Fejér* monotone sequences to be Fejér* monotone.
Remark 3.7.
Fejér* monotonicity property is preserved under synchronized linear contraction, that is, if the sequences $(x^{k})_{k\in\mathds{N}}$ and $(y^{k})_{k\in\mathds{N}}$ approach the target $z$ along the same direction with the same step-size sequence, then any convex combination of them is also Fejér* monotone. Formally, let $C$ be a closed convex set and $(x^{k})_{k\in\mathds{N}}$, $(y^{k})_{k\in\mathds{N}}\subset\mathcal{H}$ be sequences. Suppose that there exists $z\in C$, $K_{1},K_{2}\in\mathds{N}$, and scalars $\rho_{k}\in[0,1]$ such that
$x^{k+1}-z=\rho_{k}(x^{k}-z),\forall k\geq K_{1},\quad\text{and}\quad y^{k+1}-z=\rho_{k}(y^{k}-z),\forall k\geq K_{2}.$ (22)
Then, any sequence $(z^{k})_{k\in\mathds{N}}$ defined as the convex combination $z^{k}=\lambda x^{k}+(1-\lambda)y^{k}$, for $\lambda\in[0,1]$, is Fejér monotone for $k\geq K\coloneqq\max(K_{1},K_{2})$, and thus constitutes a Fejér* sequence. In fact, for $k\geq K$, we have
$z^{k+1}-z=\lambda\rho_{k}(x^{k}-z)+(1-\lambda)\rho_{k}(y^{k}-z)=\rho_{k}(z^{k}-z).$ (23)
Taking norms yields $\|z^{k+1}-z\|=\rho_{k}\|z^{k}-z\|\leq\|z^{k}-z\|$, for $k\geq K$.
@@
This section ends with the next proposition, where the closure of the convex hull of the underlying set plays a role. Even though a Fejér* monotone sequence with respect to $M\subset\mathcal{H}$ is not Fejér* monotone with respect to $\operatorname{cl}\left({\operatorname{conv}(M)}\right)$, one can characterize Fejér* monotonicity in terms of the convergence of scalar sequences and inner product sequences to the closure of the convex hull.
Proposition 3.8.
@@
Let $(x^{k})_{k\in\mathds{N}}\subset\mathcal{H}$ be Fejér* monotone with respect to a nonempty set $M$ in $\mathcal{H}$. Then,
@@
    for every $\bar{x}\in\operatorname{cl}\left({\operatorname{conv}(M)}\right)$ the scalar sequence $\left(\|{x^{k}-\bar{x}}\|\right)_{k\in\mathds{N}}$ converges;
@@
Proof.
Regarding item (i), recall that Proposition˜3.1(iii) implies that $(x^{k})_{k\in\mathds{N}}$ is Fejér* monotone with respect to $\operatorname{conv}(M)$. Thus, for any $w\in\operatorname{conv}(M)$, Proposition˜3.1(ii) yields that the scalar sequence $(\left\lVert x^{k}-w\right\rVert)_{k\in\mathds{N}}$ converges.
Now, take $\bar{x}\in\operatorname{cl}\left({\operatorname{conv}(M)}\right)$, and consider a sequence $(w^{\ell})_{\ell\in\mathds{N}}\subset\operatorname{conv}(M)$ such that $w^{\ell}\rightarrow\bar{x}$. Using the triangle inequality, we have, for all $\ell\in\mathds{N}$,
@@
Taking $\liminf$ and $\limsup$ with respect to $k$ in the above inequalities, we obtain
$-\left\|w^{\ell}-\bar{x}\right\|$ $\leq\liminf_{k\to\infty}\left\|x^{k}-\bar{x}\right\|-\lim_{k\to\infty}\left\|x^{k}-w^{\ell}\right\|$ (25)
$\leq\limsup_{k\to\infty}\left\lVert x^{k}-\bar{x}\right\rVert-\lim_{k\to\infty}\left\|x^{k}-w^{\ell}\right\|\leq\left\|w^{\ell}-\bar{x}\right\|,$ (26)
because, as aforementioned, item Proposition˜3.1(ii) says that, for every $\ell\in\mathds{N}$, $\lim_{k\to\infty}\left\|w^{\ell}-x^{k}\right\|$ exists. Now, as $\ell$ goes to infinity, we have $\left\|w^{\ell}-\bar{x}\right\|\to 0$, and we conclude that
@@
Now, item (i) implies that the sequences $\left(\left\|x^{k}-\bar{x}_{1}\right\|\right)_{k\in\mathds{N}}$ and $\left(\left\|x^{k}-\bar{x}_{2}\right\|\right)_{k\in\mathds{N}}$ converge. Therefore, the sequence $\left(\left\langle\bar{x}_{1}-\bar{x}_{2},x^{k}\right\rangle\right)_{k\in\mathds{N}}$ converges.
@@ 4 Fejér* and quasi-Fejér sequences
In this section we discuss the relation between Fejér* monotonicity and quasi-Fejér monotonicity (as addressed in [Combettes:2001a]). We use the following notation: $\ell_{+}$ is the set of real sequences with nonnegative entries, and $\ell^{1}$ is the set of real sequences with finite $\ell_{1}$-norm.
Definition 4.1 (quasi-Fejér monotonicity [Combettes:2001a, Def. 1.1]).
@@
    $\left\|x^{k+1}-x\right\|\leq\left\|x^{k}-x\right\|+\varepsilon_{k};$ (30)
@@
    $\left\|x^{k+1}-x\right\|^{2}\leq\left\|x^{k}-x\right\|^{2}+\varepsilon_{k};$ (31)
@@
    $\left\|x^{k+1}-x\right\|^{2}\leq\left\|x^{k}-x\right\|^{2}+\varepsilon_{k}.$ (32)
It is clear that Fejér sequences are quasi-Fejér of Type I sequences. Also, one can clearly see that Type II implies Type III. Moreover, it was established in [Combettes:2001a, Prop. 3.2] that if $M$ is bounded, then Type I implies Type II.
Quasi-Fejér of Type III sequences are the most general ones. However, due to this generality, they are hardly used for the analysis of optimization algorithms. In fact, the most common type of quasi-Fejér sequences used in the literature are quasi-Fejér of Type II sequences; see, for instance, [Alber:1998, Iusem:1994].
The next theorem shows that Fejér* monotone sequences are also quasi-Fejér of Type III.
Theorem 4.2.
Let $(x^{k})_{k\in\mathds{N}}\subset\mathcal{H}$ be a Fejér* monotone sequence with respect to a nonempty set $M$ in $\mathcal{H}$. Then, $(x^{k})_{k\in\mathds{N}}$ is a quasi-Fejér of Type III sequence with respect to $M$.
Proof.
Suppose that $(x^{k})_{k\in\mathds{N}}$ is Fejér* monotone with respect to $M$. Then, in view of Eq.˜2, for any $x\in M$, there exists $N(x)$ such that, for all $k\geq N(x)$, it holds that
$\left\|x^{k+1}-x\right\|^{2}\leq\left\|x^{k}-x\right\|^{2}.$ (33)
@@
$\left\|x^{k+1}-x\right\|>\left\|x^{k}-x\right\|,$ (34)
@@
$\varepsilon_{k}\coloneqq\begin{cases*}\left\|x^{k+1}-x\right\|^{2}-\left\|x^{k}-x\right\|^{2},&\text{if $\left\|x^{k+1}-x\right\|>\left\|x^{k}-x\right\|$},\\ 0,&\text{otherwise}.\end{cases*}$ (35)
Thus, we conclude that $(\varepsilon_{k})_{k\in\mathds{N}}\in\ell_{+}\cap\ell^{1}$ and, therefore, the sequence $(x^{k})_{k\in\mathds{N}}$ is quasi-Fejér of Type III with respect to $M$. ∎
Remark 4.3 (Fejér* versus quasi-Fejér sequences).
We note that the concepts of Fejér* and quasi-Fejér of Type III sequences clearly are not equivalent, with the latter being more general than the former. Moreover, if $M$ is a singleton, then any Fejér* monotone sequence is quasi-Fejér of Type II, using the same arguments used in the proof of Theorem˜4.2. Of course, in this case one can easily build quasi-Fejér sequences of Type II that are not Fejér*. Nevertheless, if a sequence is quasi-Fejér of Type II with respect to a general nonemtpy $M$, then it is Type II with respect to $\operatorname{cl}\left({\operatorname{conv}(M)}\right)$ (see [Combettes:2001a, Prop. 3.4]). Hence, in view of Example˜3.3, we have a Fejér* sequence that is not quasi-Fejér of Type II. Note that we can construct a non-singleton set $M$, for instance, $M\coloneqq[-1,0]\times\{0\}\subset\mathds{R}^{2}$, such that the same sequence $(x^{k})_{k\in\mathds{N}}$ of Example˜3.3 is quasi-Fejér of Type II, but not Fejér* monotone with respect to $M$. Therefore, more than not being equivalent, we have just seen that the concepts of Fejér* and quasi-Fejér of Type II do not imply each other, that is to say, these concepts are independent.
The previous theorem reveals a particular quasi-Fejér of Type III sequence that is related to optimization algorithms (as seen in Section˜2), through Fejér* monotonicity. Thus, we end our discussion by presenting some characterizations of Fejér* monotonicity in terms of the weak convergence of the underlying sequence. These statements are inspired by [Combettes:2001a] and could be derived from the results therein, since Fejér* monotonicity implies quasi-Fejér of Type III monotonicity.
From now on, for a sequence $(x^{k})_{k\in\mathds{N}}\subset\mathcal{H}$, we denote by $\mathfrak{W}\left(x^{k}\right)_{k\in\mathds{N}}$ the set of its weak cluster points and by $\mathfrak{S}\left(x^{k}\right)_{k\in\mathds{N}}$ the set of its strong cluster points. We remind that for a nonempty set $M$, the affine hull of $M$ is $\operatorname{aff}(M)\coloneqq\left\{\sum_{i=1}^{p}\lambda_{i}x_{i}\mid x_{i}\in M,\,\lambda_{i}\in\mathds{R},\,\sum_{i=1}^{p}\lambda_{i}=1,p\in\mathds{N}\right\},$ that is, the smallest affine subspace containing $M$. First, we give portrayals of Fejér* monotonicity in terms of the weak convergence of the sequence.
Proposition 4.4.
Let $\left(x^{k}\right)_{k\in\mathds{N}}$ be Fejér* sequence with respect to nonempty $M\subset\mathcal{H}$. Then,
@@
    $M\subset\left\{y\in\mathcal{H}\mid\left\langle y,w^{\prime}-w^{\prime\prime}\right\rangle=\alpha\right\};$ (36)
@@
Proof.
Item (i): the statement is valid since, due to Proposition˜3.1(i), the sequence $\left(x^{k}\right)_{k\in\mathds{N}}$ is bounded.
@@
$\left\|x^{k}-y\right\|^{2}-\|y\|^{2}=\left\|x^{k}\right\|^{2}-2\left\langle y,x^{k}\right\rangle.$ (37)
Moreover, it follows from Proposition˜3.1(ii) that $\lim_{k\to\infty}\left(\left\|x^{k}-y\right\|^{2}-\|y\|^{2}\right)$ exists.
Now, given two points $w^{\prime},w^{\prime\prime}\in\mathfrak{W}\left(x^{k}\right)_{k\in\mathds{N}}$, such that $x^{j_{k}}\rightharpoonup w^{\prime}$ and $x^{\ell_{k}}\rightharpoonup w^{\prime\prime}$, and considering Eq.˜37, we have
$\lim_{k\to\infty}\left\|x^{j_{k}}\right\|^{2}-2\left\langle y,w^{\prime}\right\rangle=\lim_{k\to\infty}\left\|x^{\ell_{k}}\right\|^{2}-2\left\langle y,w^{\prime\prime}\right\rangle,$ (38)